OpenStudy (anonymous):

Radical problem: square root of 50 minus square root of 32 + square root of 18= ??

OpenStudy (anonymous):

\[\sqrt{50}-\sqrt{32}+\sqrt{18} = \sqrt{25*2}-\sqrt{16*2}+\sqrt{9*2}= \sqrt{2}(5-4+3)=4\sqrt{2}\] will be the answer i hope...

OpenStudy (anonymous):

how did you get this @sriramkumar ? √(5−4+3)

OpenStudy (anonymous):

take sqrt(2) common frrom all the terms you see... @alfers101

OpenStudy (amorfide):

each of those square roots can be simplified. to simplify a surd you look for the highest square number that can possibly go into that number so for root 50 what is the highest square number to go into 50? 25 you multiply it by 2 this allows you to re write root 50 into root 25 x root 2  root 25 is 5 so you have 5xroot2 which is 5root2  follow the same process with your other square roots then you just subtract the coefficient of each square root and have the same root (root 2)

OpenStudy (anonymous):

thank you :D